# Not All Data are Good Labels: On the Self-supervised Labeling for Time Series Forecasting

### v2

@@ 3.3 Self-Correction with Adaptive Mask (Scam)
3.4 Spectral Norm Regularization (snr)
Note that we have omitted the gradient constraints in Eq. 3. This term is, in fact, positively correlated with $\nabla_{\theta}f(x;\theta)$ (discussed in Appendix D). This further supports our previous analysis, as we have only tested MLP models, which converge easily with lower $\nabla_{\theta}f$. However, when the predictor is replaced by a Transformer-based model, the dominant source of overfitting shifts from $\ell_{target}$ to $\ell_{pred}$. This phenomenon is not unique to our self-supervised learning paradigm. For example, (Ilbert et al., 2024) proposed sharpness-aware optimization to address overfitting in supervised settings. While gradient penalties are theoretically effective, they may not be practical for complex models like Transformers, as computing second-order derivatives can significantly hinder optimization.
A more direct approach is to regularize parameters using the sharpness metric, known as Spectral Norm Regularization (Miyato et al., 2018; Yoshida & Miyato, 2017):
@@ 3.4 Spectral Norm Regularization (snr)
where $\|\cdot\|_{2}$ is the spectral norm (the largest eigenvalue of parameter matrix) and $\gamma$ is a learnable scale factor. When applying to self-attention (SA) in Transformer-based architecture, snr significantly undermines the expressiveness of attention score matrices (entropy collapse (Zhai et al., 2023)). Hence, Ilbert et al. (2024) conclude that snr is inapplicable to SA parameters. However, we observe that linear layers — typically the embedding layer before SA and the projection layer after SA — also contribute to the overall sharpness of the loss landscape. Consequently, we propose applying snr selectively to the pre- and post-SA linear layers. In this way, snr can work with MLP-based models as they are also composed of multiple linear layers. Further empirical studies on snr are discussed in Section 4.2.2.
Table 1: Performance boost by adding Scam and snr to different backbones with better results in bold.
@@
4 Experiment and Analysis
We address three major questions for experiments:
Our main evaluation involves seven datasets: Electricity, Weather, Traffic, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), all of which are well-established TSF benchmarks and publicly available (Wu et al., 2021). We also test the proposals using four PeMS datasets of a larger scale, as reported in Appendix 6. The predictor (i.e., the backbone model integrated with Scam) covers representative TSF models, including both MLP-based and Transformer-based architectures. Mlp (Li et al., 2023) is a vanilla 2-layer baseline equipped with RevIN (Kim et al., 2021) while cycleNet (Lin et al., 2024b) is a SOTA MLP-based model explicitly capturing cyclic trend in time series. PatchTST (Nie et al., 2022) and iTransformer (Liu et al., 2024b) are Transformer-based models, representing channel-independent and channel-dependent methods, respectively. Following previous settings (Zhou et al., 2021; Wu et al., 2021, 2023) for direct and fair comparison, we set prediction length $H\in\{96,192,336,720\}$ and look-back length to 96 for all datasets. We provide dataset descriptions, implementation details, and reproduction instructions in Appendix F.
4.1 Main Experiment (Q1)
Table 1 demonstrates consistent performance improvements in all major backbones when Scam and snr are incorporated in the self-supervised-learning paradigm. The full results with detailed breakdowns by prediction lengths are provided in Appendix 6. These gains are particularly notable on ETT datasets, which are known for their noisy nature and relatively small size. Notably, Transformer-based models like PatchTST and iTrans, which typically underperform compared to lightweight models Mlp and cycleNet on these datasets, show significant enhancements in generalization with Scam. On the Weather dataset, the boost is more modest, likely due to the intrinsic chaotic nature of atmospheric data.
Regarding Q1, our method demonstrates general effectiveness across various backbones and datasets. A well-known discrepancy between MLP-based and Transformer-based models is their dataset preferences: Transformer-based methods excel on large, regular datasets, while MLP-based methods perform better on noisy datasets. Scam helps bridge this gap by enabling Transformer-based models to perform competitively on traditionally challenging datasets and enhancing the robustness of MLP-based models.
@@ 4.2 Ablation Study (Q2)
4.2.1 Contributions of Scam and snr
Table 2: Ablation study on Scam and snr for cycleNet.
@@ 4.2.1 Contributions of Scam and snr
Table 3: Ablation study on Scam and snr for iTrans.
@@
To answer Q2, we present the performance gains from Scam and snr through an ablation study (Tables 3 and 3), using iTrans (Transformer-based) and cycleNet (MLP-based) as representatives (the SOTA one in their category).
snr, a practical alternative to the gradient penalty in Eq. 9, consistently enhances performance across backbones. However, iTrans, being more prone to overfitting on small datasets (discussed in (Ilbert et al., 2024)), benefits less from snr compared to cycleNet, likely due to insufficient data for generalization in higher-complexity architectures. With Scam, both models exhibit significant improvements, leveraging the expanded effective training data.
@@
4.2.2 In-depth Examination on snr
We further verify the effectiveness of snr by analyzing the loss landscape. Transformer-based models, as discussed, are more prone to overfitting, particularly on noisy datasets. This weakens our adaptive mask’s ability to discard overfitted components, as $\ell_{pred}$ — the predictor’s training loss — also contributes to overfitting. Gradient penalty offers a solution to this issue. For TSF models specifically, where both inputs and outputs are time series with abrupt distribution shifts, parameter robustness is critical for mitigating overfitting. Previous works (Yoshida & Miyato, 2017; Miyato et al., 2018) have shown that regulating parameters via their spectral norm enhances stability against input perturbations, which is particularly beneficial in TSF.
Figure 8: iTrans as a TST example
In Figure 9, the bars represent the sharpness of different components in iTrans, divided into three parts: embedding, encoder, and projector (see Figure 8). The embedding and projector are linear layers, while the encoder comprises channel-wise attention blocks. This architectural pattern is common among Time Series Transformers (TSTs), with the encoder being the primary focus and less emphasis placed on pre- or post-encoder linear layers. Without snr, the input linear layer (embedding) exhibits the highest sharpness, indicating it is the most overfitted component (see red bars). This suggests that overfitting in Transformers originates not only from self-attention mechanisms but also from the linear layers. By applying snr to the pre- and post-encoder linear layers (Section 3.4), we observe smoother loss landscapes, validating snr’s ability to reduce sharpness effectively.
Figure 9: Sharpness of difference components in iTrans.
Figure 10: Empirical results on different variants of snr.
@@ 4.3 Scam: A Multiple Instance Learning View (Q3)
Multiple Instance Learning (MIL) is a classical weakly-supervised binary classification problem where typically a bag of instances is labeled positive if at least one instance is positive. (Early et al., 2024; Chen et al., 2024b) extend MIL to Time Series Classification by treating an input window as a bag of instances (time points), enabling a model to predict based on instance-level classifications that are more interpretable.
@@
This study only reveals a part of Scam’s self-supervision effectiveness, which is further explored in Appendix D.
@@ 5 Related Work
We discuss related techniques from meta-learning and self-supervised learning perspectives, with an inventory of TSF models in Appendix E.
Meta-Learning for Time Series. Meta-Learning, by definition, seeks to perform in a learn-to-learn paradigm. Generally speaking, meta-learning includes optimization-based, model-based, and metric-based methods. Optimization-based methods target optimal initial parameters (Finn et al., 2017), often involving a two-loop optimization process (Oreshkin et al., 2021; Woo et al., 2022). Model-based methods aim to determine suitable models, typically from an ensemble, based on predefined tasks (Lines et al., 2018; Middlehurst et al., 2021) or activation states (Abdallah et al., 2022). Metric-based methods (Du et al., 2021; Woo et al., 2022) learn a metric function that provides more expressive measurements of distances between data samples, commonly comparing training and validation samples.
Our method Scam aligns with the boarder scope of meta-learning. Specifically, the grid search (Algorithm 1) follows the two-loop structure similar to optimization-based methods. However, it diverges by focusing on dataset space rather than parameter space. The goal of the grid search is not to solve the optimization problem directly but to leverage the trajectory of optimization. Additionally, the final mask form of Scam in essence provides a more accurate metric tailored for a supervised-learning setting. While DeepTime (Woo et al., 2022) and AdaRNN (Du et al., 2021) share a related idea, there are key differences. DeepTime focuses more on a time-index forecast paradigm, whereas AdaRNN applies metric learning to hidden states. Both works learn metrics on the sample level (a window of time series), whereas ours focuses on the instance level (individual data points of time series).
Self-supervised Learning for Time Series. Self-supervised learning trains models without relying on manually labeled data by using auxiliary tasks like generation, contrast, and reconstruction to learn expressive representation or create pseudo labels. In the realm of time series, this approach is discussed more in Time Series Classification (TSC) (Jawed et al., 2020; Xi et al., 2022; Liu et al., 2024c). Recent works (Early et al., 2024; Chen et al., 2024b) present a novel perspective that instances in time series/segments can have multiple labels. They propose corresponding weakly-supervised-learning methods that significantly improve both performance and interpretability.
As manual labeling is usually not required, TSF is treated as a generation task by self-supervised methods. Recent works focus on the use of TSF as an auxiliary task to learn universal representations that improve performances of other tasks (Nie et al., 2022; Senane et al., 2024; Liu & Chen, 2024). This paradigm shows the potential to scale time series models to levels comparable to large language models.
In this work, we integrate both perspectives by employing an auxiliary reconstruction task, commonly used in TSC, to enhance the performance of TSF. The pseudo labels, often discussed in TSC, are derived from existing, manual ones, while ours are created in a self-supervised paradigm.
@@ 6 Conclusion and Future Work
This paper presents a self-supervised approach Scam that enhances TSF models by selectively replacing overfitted components with pseudo labels derived from intermediate reconstructions. Combined with Spectral Norm Regularization applied to the linear layers of the backbone, Scam improves generalization and robustness across TSF models. Future work will explore extending Scam to tasks such as time series outlier detection and error correction.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Grid Search Algorithm in Initial Case
@@
5for $i\leftarrow 0$ to $N$ do
       $\triangleright$ $N$ for number of candidates
@@ 25      Backpropagate $\ell_{rec}$ and update $\phi$ using $Opt_{\phi}$;
Appendix B Further Discussion on Gradient Constraint
@@
[…]
@@
$\quad+\beta\|\nabla_{\theta}f(x;\theta)\|,$
For readers who are familiar with Reinforcement Learning, this derivation resembles the transfer from a constrained optimization to a penalized one (e.g., from TRPO (Schulman, 2015) to PPO (Schulman et al., 2017)). In brief, while the penalized form is theoretically equivalent to the constrained form, it is challenging to choose a fixed $\beta$ that works universally across all datasets or even within a single dataset (because intrinsic characteristics can vary over time). Thus, a more general form of constraint is required to better serve the penalty, similar to the concept of gradient clipping in PPO.
@@
$\|\sigma^{\prime}_{1}W(x1-x2)\|\leq C(W,b)\|x_{1}-x_{2}\|,$ (11)
When the gradient of activation function $\sigma$ has an upper bound (as is often the case for common activation functions like ReLU, Sigmoid, etc.), the Lipchitz condition holds as long as
@@
However, in practice, snr have limitations when applied to the parameter matrix in self-attention mechanisms. This phenomenon is termed entropy catastrophe as discussed by (Ilbert et al., 2024). In this paper, by analyzing the sharpness of different components in the predictor model, we propose to use pre-snr and post-snr combined, which specifically normalizes the first and last linear layer in the TSF models (see Section 3.4).
Appendix C Implementation Details of $g(\cdot;\phi)$
As introduced in Section 2.2, we propose a simple enough reconstruction network $g(\cdot;\phi)$ that serves our objective. Despite its simplicity, the architecture incorporates some special designs that enhance its performance. Specifically, these include the conv-concat layer and the point-wise FFN, which are detailed in Appendix C.1 and Appendix C.2, respectively.
C.1 Conv-concat Layer
@@
[…]
@@
Effective Receptive Field. As what was proposed by (Luo & Wang, 2024), the Effective Receptive Field (ERF) is a reasonable consideration for designing convolution-based architectures. To evaluate the ERF of our conv-concat layer, we input an impulse function and visualize the resulting ERF, as shown in Figure 13. The visualization demonstrates that, without requiring an extra-large convolution kernel, our proposed method achieves a near-global ERF. This is made possible by combining the outputs from different layers, each capturing distinct frequency patterns.
@@
C.2 Point-wise FFN
We employ a point-wise FFN as a parameter-sharing module to decode the outputs from the convolution layer. The FFN is essentially a two-layer MLP that resembles the common design of a linear projector in predictor models, as mentioned in Section 4.2.2.
@@ C.2 Point-wise FFN
  • Patch-dependent Design: PatchTST (Nie et al., 2022) adopts a patch-dependent linear projector (shown in Figure 14(a)) that first flattens all features within patches and utilizes an extra-large weight matrix of shape $Ld\times d$, where $L$ is the sequence length and $d$ is the dimension of the latent embedding.
  • Patch-independent Design: (Lee et al., 2023) proposes that the necessity for patch-dependent designs depends on the specific task. For instance, tasks like forecasting may require patch-dependent projectors, while tasks like contrastive learning might favor a patch-independent design (shown in Figure 14(b)).
@@
(a) patch-dependent linear projector
(b) patch-independent linear projector
(c) point-wise linear projector
@@
Appendix D Further Analysis
D.1 Distribution of Reconstructed Datasets in Grid Search

### v3

@@ 3.4 Spectral Norm Regularization (snr)
A more direct approach is to regularize parameters using the sharpness metric, known as Spectral Norm Regularization [26, 46]:
@@
where $\|\cdot\|_{2}$ is the spectral norm (the largest eigenvalue of parameter matrix) and $\gamma$ is a learnable scale factor. When applying to self-attention (SA) in Transformer-based architecture, snr significantly undermines the expressiveness of attention score matrices (entropy collapse [49]). Hence, Ilbert et al. [10] conclude that snr is inapplicable to SA parameters. However, we observe that linear layers — typically the embedding layer before SA and the projection layer after SA — also contribute to the overall sharpness of the loss landscape. Consequently, we propose applying snr selectively to the pre- and post-SA linear layers. In this way, snr can work with MLP-based models as they are also composed of multiple linear layers. Further empirical studies on snr are discussed in Section E.3.
4 Experiment and Analysis
We address three major questions for experiments:
Our main evaluation involves seven datasets: Electricity, Weather, Traffic, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), all of which are well-established TSF benchmarks and publicly available [40]. We also test the proposals using four PeMS datasets of a larger scale, as reported in Appendix 7. The predictor (i.e., the backbone model integrated with Scam) covers representative TSF models, including both MLP-based and Transformer-based architectures. Mlp [15] is a vanilla 2-layer baseline equipped with RevIN [12] while cycleNet [17] is a SOTA MLP-based model explicitly capturing cyclic trend in time series. PatchTST [27] and iTransformer [21] are Transformer-based models, representing channel-independent and channel-dependent methods, respectively. Following previous settings [53, 40, 41] for direct and fair comparison, we set prediction length $H\in\{96,192,336,720\}$ and look-back length to 96 for all datasets. We provide dataset descriptions, implementation details, and reproduction instructions in Appendix G.
4.1 Main Experiment (Q1)
Table 1: Performance boost by adding Scam and snr to different backbones. Better results are in bold.
@@ 4.1 Main Experiment (Q1)
Table 1 demonstrates consistent performance improvements in all backbones across all datasets when Scam and snr are incorporated in the self-supervised-learning paradigm. The full results with detailed breakdowns by prediction lengths are provided in Appendix 7. These gains are particularly notable on ETT datasets, which are known for their noisy nature and relatively small size. Notably, Transformer-based models like PatchTST and iTrans, which typically underperform compared to lightweight models Mlp and cycleNet on these datasets, show significant enhancements in generalization with Scam. On the Weather dataset, the boost is more modest, likely due to the intrinsic chaotic nature of atmospheric data.
Regarding Q1, our method demonstrates general effectiveness across various backbones and datasets. A well-known discrepancy between MLP-based and Transformer-based models is their dataset preferences: Transformer-based methods excel on large, regular datasets, while MLP-based methods perform better on noisy datasets [30, 34]. Scam helps bridge this gap by enabling Transformer-based models to perform competitively on traditionally challenging datasets and enhancing the robustness of MLP-based models.
@@ 4.2 Ablation Study (Q2)
Table 2: Ablation study for cycleNet.
@@
Table 3: Ablation study for iTrans.
@@
To answer Q2, we present the performance gains from Scam and snr through an ablation study (Tables 3 and 3  [^2] The results in both tables may have discrepancy in baseline results from Table 1. This is because the Table 1 has baseline results from the original papers, and Table 3 and 3 are from our runs for module modification. See Sec G.3 for further explanations.), using iTrans (Transformer-based) and cycleNet (MLP-based) as representatives (the SOTA one in their category).
snr, a practical alternative to the gradient penalty in Eq. 8, consistently enhances performance across backbones. However, iTrans, being more prone to overfitting on small datasets (discussed in [10]), benefits less from snr compared to cycleNet, likely due to insufficient data for generalization in higher-complexity architectures. With Scam, both models exhibit significant improvements, leveraging the expanded effective training data.
@@ 4.3 Scam: A Multiple Instance Learning View (Q3)
Multiple Instance Learning (MIL [25, 3]) is a classical weakly-supervised binary classification problem where typically a bag of instances is labeled positive if at least one instance is positive. [8, 5] extend MIL to Time Series Classification by treating an input window as a bag of instances (time points), enabling a model to predict based on instance-level classifications that are more interpretable.
@@
This study only reveals a part of Scam’s self-supervision effectiveness, which is further explored in Appendix E.
@@ 5 Related Work
We discuss related techniques from meta-learning and self-supervised learning perspectives, with an inventory of TSF models in Appendix F.
Meta-Learning for Time Series. Meta-Learning, by definition, seeks to perform in a learn-to-learn paradigm. Generally speaking, meta-learning for TSF includes optimization-based and metric-based methods. Optimization-based methods target optimal initial parameters [9], often involving a two-loop optimization process [28, 38]. Metric-based methods [7, 38] learn a metric function that provides more expressive measurements of distances between data samples, commonly comparing training and validation samples.
Our method Scam aligns with the broader scope of meta-learning. Specifically, the grid search (Section 3.1, Appendix B) follows the two-loop structure similar to optimization-based methods. However, it diverges by focusing on dataset space rather than parameter space. Additionally, the final mask form of Scam in essence provides a more accurate metric tailored for a supervised-learning setting. Previous works [38, 7] learn metrics on the sample level (a window of time series), whereas ours focuses on the instance level (individual data points of time series).
Self-supervised Learning for Time Series. Self-supervised learning trains models without relying on manually labeled data by using auxiliary tasks like generation, contrast, and reconstruction to learn expressive representation or create pseudo labels. In the realm of time series, this approach is discussed more in Time Series Classification (TSC) [11, 42, 22]. Recent works [8, 5] present a novel perspective that instances in time series/segments can have multiple labels. They propose corresponding weakly-supervised-learning methods that significantly improve both performance and interpretability.
In this work, we integrate the multiple-label perspective by employing an auxiliary reconstruction task, commonly used in TSC, to enhance the performance of TSF. The pseudo labels, often discussed in TSC, are derived from existing, manual ones, while ours are created in a self-supervised paradigm.
@@ 6 Conclusion and Future Work
This paper presents a self-supervised approach Scam that enhances TSF models by selectively replacing overfitted components with pseudo labels derived from intermediate reconstructions. When combined with Spectral Norm Regularization, Scam improves generalization and robustness across TSF models. Future work will explore extending Scam to tasks such as time series outlier detection and error correction.
7 Acknowledgment
This work was supported by the National Natural Science Foundation of China (Nos. 62402420, 62402414), Fundamental Research Funds for Central Universities (Grant No. 226202400109), the Zhejiang Provincial Natural Science Foundation (No. LD24F020015), and the National Regional Innovation and Development Joint Fund (No. U24A20254). Huan Li is the corresponding author of this paper.
Appendix A Limitations
Our work constrains the discussion for TSF models in MLP- and Transformer-based ones, excluding other potentially feasible solutions [23, 1]. We omit these newly explored architectures since they have not outperformed the baseline methods we have adopted under the same budget. Also, we have not verified our method on LLM-based models (or models of an LLM-level scale) [39, 2] due to the excessive costs required.
Appendix B Grid Search Algorithm in Initial Case
@@
5for $i\leftarrow 0$ to NN do
 $\triangleright$ NN for number of candidates
@@ 25  Backpropagate $\ell_{rec}$ and update $\phi$ using $Opt_{\phi}$;
Appendix C Further Discussion on Gradient Constraint
@@
$\approx\begin{bmatrix}\nabla_{\phi}g&-\nabla_{\theta}f\end{bmatrix},\quad\text{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}because $|\tilde{y}-y|\ll|\tilde{y}-\hat{y}|$}.}$
@@
$\quad+\beta\|\nabla_{\theta}f(x;\theta)\|.$
For readers who are familiar with Reinforcement Learning, this derivation resembles the transfer from a constrained optimization to a penalized one (e.g., from TRPO [32] to PPO [33]). In brief, while the penalized form is theoretically equivalent to the constrained form, it is challenging to choose a fixed $\beta$ that works universally across all datasets or even within a single dataset (because intrinsic characteristics can vary over time). Thus, a more general form of constraint is required to better serve the penalty, similar to the concept of gradient clipping in PPO.
@@
$\quad\ \left\|\sigma\left(Wx_{1}\right)-\sigma\left(Wx_{2}\right)\right\|$ (10)
$\approx\left\|\left(\sigma(x_{0})+\sigma^{\prime}(x_{0})\left(Wx_{1}-x_{0}\right)\right)-\left(\sigma(x_{0})+\sigma^{\prime}(x_{0})\left(Wx_{2}-x_{0}\right)\right)\right\|,\quad\text{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}by Taylor expansion}}$
$=\left\|\sigma^{\prime}(x_{0})\left(W\left(x_{1}-x_{2}\right)\right)\right\|$
$\leq\sigma^{\prime}_{max}\|W\|\|x_{1}-x-2\|$
$\leq C(W)\|x_{1}-x_{2}\|.$
When the derivative of activation function $\sigma$ has an upper bound $\sigma^{\prime}_{max}$(as is often the case for common activation functions like ReLU, Sigmoid, etc.), the Lipchitz condition holds as long as
@@
However, in practice, snr have limitations when applied to the parameter matrix in self-attention mechanisms. This phenomenon is termed entropy catastrophe as discussed by [10]. In this paper, by analyzing the sharpness of different components in the predictor model, we propose to use pre-snr and post-snr combined, which specifically normalizes the first and last linear layer in the TSF models (see Section 3.4).
Appendix D Implementation Details of $g(\cdot;\phi)$
As introduced in Section 2.2, we propose a simple enough reconstruction network $g(\cdot;\phi)$ that serves our objective. Despite its simplicity, the architecture incorporates some special designs that enhance its performance. Specifically, these include the conv-concat layer and the point-wise FFN, which are detailed in Appendix D.1 and Appendix D.2, respectively.
D.1 Conv-concat Layer
@@
Specifically, we set $\text{kernel size}=3$, $\text{stride}=2$, $\text{padding}=1$, and the number of kernels doubled for each subsequent layer. Using this setup, as illustrated in Figure 9, we ensure that the number of features remains invariant across different layers, with only the shape of the features changing. Now we can fuse the outputs from different convolution layers together by flattening/unfolding the features to the original shape of ($L\times 1$). Again, considering the effectiveness of point-wise FFN presented in Appendix D.2, we expect the concatenated features to be near-equally arranged along the temporal dimension to preserve the sequential relationships in the embedding. To achieve this, we first transpose the features and then unfold them. This practice can ensure a Receptive Field of $(2^{l+1}-1)$ wide for each embedding, where $l$ is the total number of convolution layers.
@@
Effective Receptive Field. As what was proposed by [23], the Effective Receptive Field (ERF) is a reasonable consideration for designing convolution-based architectures. To evaluate the ERF of our conv-concat layer, we input an impulse function and visualize the resulting ERF, as shown in Figure 10. The visualization demonstrates that, without requiring an extra-large convolution kernel, our proposed method achieves a near-global ERF. This is made possible by combining the outputs from different layers, each capturing distinct frequency patterns.
@@
D.2 Point-wise FFN
We employ a point-wise FFN as a parameter-sharing module to decode the outputs from the convolution layer. The FFN is essentially a two-layer MLP that resembles the common design of a linear projector in predictor models, as mentioned in Section E.3.
@@ D.2 Point-wise FFN
  • Patch-dependent Design: PatchTST [27] adopts a patch-dependent linear projector (shown in Figure 11(a)) that first flattens all features within patches and utilizes an extra-large weight matrix of shape $Ld\times d$, where $L$ is the sequence length and $d$ is the dimension of the latent embedding.
  • Patch-independent Design: [14] proposes that the necessity for patch-dependent designs depends on the specific task. For instance, tasks like forecasting may require patch-dependent projectors, while tasks like contrastive learning might favor a patch-independent design (shown in Figure 11(b)).
@@
(a) patch-dependent
(b) patch-independent
(c) point-wise
@@
Appendix E Further Analysis
E.1 Distribution of Reconstructed Datasets in Grid Search